# Quantum typicality approach to energy flow between two spin-chain domains at different temperatures

### v1

@@ I Introduction
Over the last decades, the broad field of quantum many-body systems out of equilibrium has crystallized as one of the central topics of modern physics, both experimentally and theoretically [1, 2, 3, 4, 5], ranging from fundamental questions in statistical mechanics to applied questions in material science. The understanding of nonequilibrium physics has developed remarkably over the last years due to experimental advances [1], new theoretical concepts such as the typicality of pure states [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] and eigenstate thermalization [31, 32, 33], and the development of sophisticated numerical techniques [34, 35, 36, 37, 38, 39].
@@
One paradigmatic example of a nonequilibrium process is transport [40], a natural phenomenon in systems with one or more conserved quantities. Additionally, transport appears in both open and closed systems. In an open system scenario, transport processes are usually induced by coupling the system at its boundaries to external baths at different temperatures or chemical potentials. Then, in the long-time limit, a nonequilibrium steady state usually emerges with a characteristic density profile and a constant current [41, 42, 43]. Such a situation is often modeled by a quantum master equation of Lindblad form [44], which enables the application of sophisticated numerical techniques.
@@
In a closed system scenario, a widely used approach for studying transport is linear response theory, where the Kubo formula in terms of autocorrelation functions plays a crucial role [45]. This theory predicts the behavior of a system near equilibrium, and it can be formulated either in frequency and momentum space or in the space and time domain. Another approach for studying transport in closed systems is a quantum quench [46, 47, 48]. Here, a quantum system is prepared in a particular initial state, typically an eigenstate of some pre-quench Hamiltonian. Then, the system undergoes a sudden change and is described by a new Hamiltonian, and the initial state evolves unitarily according to this post-quench Hamiltonian. Quantum quenches can be studied in bipartite systems [49, 50], where the total system is composed of two subsystems and initially both subsystems are prepared in a state according to their respective Hamiltonians. Then, the subsystems are brought into contact together, and the total state evolves under the Hamiltonian of the total system.
@@
Numerically, transport in open and closed systems has been investigated using various methods, including exact and Lanczos diagonalization [34, 35, 51], simulations based on matrix product states [52, 53, 54, 42, 38, 55, 56], and Monte-Carlo techniques [57, 58, 37, 39], to name a few. Additionally, the method of dynamical quantum typicality [26, 28] is a multifaceted method for investigating transport in closed quantum systems. The basic idea of dynamical quantum typicality is that one can imitate the expectation value of an entire ensemble by an expectation value of a single pure state, drawn at random from a high-dimensional Hilbert space. Dynamical quantum typicality has been used for linear-response functions, mostly at high temperatures [40]. For such temperatures, the statistical error is smaller than at low temperatures. Still, dynamical quantum typicality is a priori neither restricted to linear-response functions nor high temperatures.
@@
In this work, we demonstrate the usefulness of dynamical quantum typicality in a specific physical situation, which differs from most previous applications in two ways. First, we apply it to energy flow in bipartite systems. Second, we also explore low temperatures. We do so for various spin-1/2 chains, which all feature ballistic energy transport: the XX chain, the critical transverse-field Ising chain, and the XXZ chain. We compare our numerical results to existing analytical results from conformal field theory (CFT) [59, 50] and generalized hydrodynamics (GHD) [49, 60, 61]. We find convincing agreement in all considered subsystem configurations of our bipartite setup, as well as for all examined temperatures. This agreement is not only found for the contact current but also for all local currents and densities in the steady state.
@@
Our paper is structured as follows: In Sec. II, we define the bipartite setup used throughout the paper and the energy current between the two subsystems. After that, we introduce the different spin-1/21/21 / 2 chains in Sec. III, which we use for our subsystem configurations. Next, in Sec. IV, we discuss our quantum typicality approach to the energy current and briefly review the CFT results. Thereafter, we present our numerical results and compare them to analytical results in Sec. V. We conclude in Sec. VI and give directions for further research.
@@ II Setup
In our paper, we are interested in the heat current from the left half into the right one. This heat current is given by [62]
@@
Figure 2: Time evolution of the energy current for temperatures $T_{L}/J=1/6,T_{R}/J=1/8$, total system size $L=28$ for the coupling of (a) two XX chains, (b) two critical Ising chains and (c) one XX chain to a critical Ising chain. Numerical data for two different pure random initial states ($N$=1) and for the average over many initial states (N=100) are compared to the CFT value.
@@ III Models
First, we introduce the spin-1/21/21 / 2 XXZ model. For this model, the local Hamiltonians are given by [40]
@@
where $S_{r}^{i}\ (i=x,y,z)$ are the spin-1/21/21 / 2 operators at site $r$, $J>0$ is the antiferromagnetic exchange coupling constant, and $\Delta$ is the anisotropy in $z$ direction. Using the Jordan-Wigner transformation [63], the spin-1/21/21 / 2 XXZ chain can be mapped onto a chain of spinless fermions. After this transformation, the local Hamiltonians in Eq. (6) take on the form
[…]
@@
Figure 3: Similar data as the one in Fig. 2 but now for different total system sizes $L$. Here, the numerical data are again averaged over $N=100$ pure random initial states. For larger system sizes $L$, finite size effects are visible for later times $tJ$.
The second model we consider is the special case of the spin-1/21/21 / 2 XXZ chain with $\Delta=0$, i.e., the XX chain. The local Hamiltonian reads
@@
where the constant $h$ denotes the strength of the transverse magnetic field. Due to the Jordan-Wigner transformation, the transverse-field Ising model is, up to a constant, equivalent to the Kitaev chain [64, 65, 66, 67],
@@
In our paper, we set $J_{C}=1$ for all configurations. All chosen coupling Hamiltonians ensure that energy can flow between both subsystems. Before we proceed to our numerical results for the energy current in the various subsystem configurations, we introduce our dynamical typicality approach. We also briefly discuss the result obtained within CFT [59].
@@ IV.1 Dynamical typicality
We now turn to a discussion of dynamical quantum typicality (DQT) [28, 26], which is the key concept used in our numerical approach. We first focus on the homogeneous case of a single Hamiltonian $H$ and a single temperature $T$, and afterwards come to the bipartite case with two $H_{L}$, $H_{R}$ and two $T_{L}$, $T_{R}$.
The basic idea of DQT [10, 13, 14] is that one can approximate the expectation value of an ensemble density matrix,
@@
where $O$ is a local operator. If $\rho$ was the microcanonical ensemble and $|\psi\rangle$ an energy eigenstate, this idea would be identical to the well-known eigenstate thermalization hypothesis (ETH) [31, 32, 33]. In contrast to ETH, however, DQT is no assumption and relies on a particular construction of the pure state $|\psi\rangle$. Specifically, for a given ensemble density matrix $\rho$, this construction reads
@@
which is the action of the square-root of $\rho$ on a so-called Haar-random pure state $|\Phi\rangle$ [16]. This particular pure state $|\Phi\rangle$ does neither depend on $\rho$ nor $H$. In an arbitrarily chosen orthogonal basis, it is given by
@@
where $D$ is the dimension of the Hilbert space and $c_{m}$ are complex coefficients. Both, the real and imaginary part of these coefficients are drawn at random according to a Gaussian probability distribution with zero mean. For such a $|\Phi\rangle$ and the resulting $|\psi\rangle$, DQT then leads to the approximation [15, 16]
@@
where the statistical error $\varepsilon$ vanishes on average. More importantly, however, the standard deviation is bounded from above,
$\sigma(\varepsilon)<b={\cal O}\left(\frac{1}{\sqrt{D_{\text{eff}}}}\right)\,.$ (19)
Here, $D_{\text{eff}}$ is the so-called effective dimension. In the case of the canonical ensemble $\rho\propto\exp(-\beta H)$, which is most relevant in the context of our work, the effective dimension is given by the partition function
@@
with the ground-state energy $E_{0}$. Thus, $D_{\text{eff}}$ counts the number of thermally occupied energy eigenstates. In the high-temperature limit $T\to\infty$, $D_{\text{eff}}=D=2^{L}$ and hence the statistical error $\varepsilon$ decreases exponentially fast with system size $L$. In the zero-temperature limit $T\to 0$, however, $D_{\text{eff}}=1$ and thus the statistical error $\varepsilon$ can be in principle large. Still, one has to take into account that the bound in Eq. (19) is not tight. In fact, at $T=0$, it turns out that the statistical error becomes $\varepsilon=0$. This fact is a kind of trivial consequence for a Haar-random pure state $|\Phi\rangle$: It is also random in the energy eigenbasis and the coefficient $c_{0}$ for the ground state |0|0$\rangle$| 0 ⟩ might be small but likely nonzero. At $T=0$, only this coefficient survives and one has $|\psi\rangle=|0\rangle$ [68].
@@
if not stated otherwise. In our case, the local observable $O$ is mostly the contact current $J_{0}$, but we also consider currents $J_{r}$ at other sites $r$ as well as the local energy densities $h_{r}$.
In our numerical simulations, we calculate the action of the imaginary-time operators $e^{-\beta_{L}H_{L}/2}$ and $e^{-\beta_{R}H_{R}/2}$ on the Haar-random pure states $|\Phi_{L}\rangle$ and $|\Phi_{R}\rangle$ with a fourth-order Runge-Kutta scheme, where we choose a very small step $\delta\beta J=0.0001$. This scheme is also used for the action of the real-time operator $e^{-iHt}$ on the pure state $|\psi\rangle$, where we choose a larger but still small step $\delta tJ=0.01$. By doing so and also using sparse-matrix techniques, we can treat system sizes outside the range of standard exact diagonalization. Alternatively, other propagation methods, such as Trotter decompositions [69, 70, 71] or Chebyshev polynomial expansions [72, 73], can be used as well.
@@ IV.2 Conformal field theory
Figure 5: Steady-state profile for (a) the local density $\langle h_{r}\rangle$ and (b) the local current $\langle J_{r}\rangle$ for the coupling of two XXZ chains with the same anisotropies $\Delta$, $T_{L}/J=1$,$T_{R}/J=1/2$, and total system size $L=24$. The numerical data are compared to the results from GHD in Ref. [49]. Here, the numerical data are shown for both, a random product state (open symbols) and a random state in the entire Hilbert space (closed symbols).
The energy transport in a bipartite CFT setup has been studied by Bernard and Doyon [59]. Specifically they considered a CFT separated into two halves prepared at different temperatures. These halves were then glued together and the non-equilibrium steady state emerging at late times was studied. In particular, for the energy current they obtained the universal formula
@@
where $c$ denotes the central charge of the CFT. Given that CFTs generically describe the universal low-energy behavior of certain lattice models, the result (24) is expected to hold only in the low-temperature regime. The central charge $c$ is a specific characteristic of the CFT [74], physically it can be interpreted as the number of degrees of freedom, which here contribute to the energy transport through the system.
The setup leading to the universal CFT result (24) possesses translation invariance after the gluing quench. As an extension, Ref. [50] studied the coupling of an XX chain to a critical Ising chain, i.e., our setup (iii) above. Using both the non-equilibrium Green function formalism and the density matrix renormalization group method, the energy current was found to behave as
@@
This result can be viewed as a bottleneck effect, where the subsystem with the lower number of degrees of freedom as reflected by the lower central charge limits the energy transport. The result (25) is also consistent with a purely field theoretic analysis by Bernard et al. [75], where, however, the precise link to a microscopic setup remains elusive.

### v2

@@ I Introduction
Over the last decades, the broad field of quantum many-body systems out of equilibrium has crystallized as one of the central topics of modern physics, both experimentally and theoretically Bloch et al. [2008], Polkovnikov et al. [2011], Eisert et al. [2015], D’Alessio et al. [2016], Abanin et al. [2019], ranging from fundamental questions in statistical mechanics to applied questions in material science. The understanding of nonequilibrium physics has developed remarkably over the last years due to experimental advances Bloch et al. [2008], new theoretical concepts such as the typicality of pure states Alben et al. [1975], De Raedt and de Vries [1989], Jaklič and Prelovšek [1994], Hams and De Raedt [2000], Gemmer and Mahler [2003], Iitaka and Ebisuzaki [2003], Gemmer et al. [2004], Goldstein et al. [2006], Popescu et al. [2006], Reimann [2007], Bartsch and Gemmer [2009], Sugiura and Shimizu [2012, 2013], Elsayed and Fine [2013], Steinigeweg et al. [2014], Monnai and Sugita [2014], Endo et al. [2018], Richter and Steinigeweg [2019], Wietek et al. [2019], Rousochatzakis et al. [2019], Heitmann et al. [2020], Schnack et al. [2020], Jin et al. [2021], Reimann [2018], Mitrić [2025] and eigenstate thermalization Deutsch [1991], Srednicki [1994], Rigol et al. [2008], and the development of sophisticated numerical techniques Jaklič and Prelovšek [2000], Long et al. [2003], Schollwöck [2005], Grossjohann and Brenig [2010], Schollwöck [2011], De Raedt et al. [2017].
@@
One paradigmatic example of a nonequilibrium process is transport Bertini et al. [2021], a natural phenomenon in systems with one or more conserved quantities. Additionally, transport appears in both open and closed systems. In an open system scenario, transport processes are usually induced by coupling the system at its boundaries to external baths at different temperatures or chemical potentials. Then, in the long-time limit, a nonequilibrium steady state usually emerges with a characteristic density profile and a constant current Michel et al. [2003], Prosen and Žnidarič [2009], Žnidarič [2011], Xu et al. [2023]. Such a situation is often modeled by a quantum master equation of Lindblad form Breuer and Petruccione [2007], which enables the application of sophisticated numerical techniques.
@@
In a closed system scenario, a widely used approach for studying transport is linear response theory, where the Kubo formula in terms of autocorrelation functions plays a crucial role Kubo et al. [1991]. This theory predicts the behavior of a system near equilibrium, and it can be formulated either in frequency and momentum space or in the space and time domain. Another approach for studying transport in closed systems is a quantum quench Calabrese and Cardy [2006], Mitra [2018], Richter et al. [2018]. Here, a quantum system is prepared in a particular initial state, typically an eigenstate of some pre-quench Hamiltonian. Then, the system undergoes a sudden change and is described by a new Hamiltonian, and the initial state evolves unitarily according to this post-quench Hamiltonian. Quantum quenches can be studied in bipartite systems Bertini et al. [2016], Fischer et al. [2020], where the total system is composed of two subsystems and initially both subsystems are prepared in a state according to their respective Hamiltonians. Then, the subsystems are brought into contact together, and the total state evolves under the Hamiltonian of the total system.
@@
Numerically, transport in open and closed systems has been investigated using various methods, including exact and Lanczos diagonalization Jaklič and Prelovšek [2000], Long et al. [2003], Karrasch et al. [2013], simulations based on matrix product states Vidal [2004], Zwolak and Vidal [2004], Verstraete et al. [2008], Prosen and Žnidarič [2009], Schollwöck [2011], Weimer et al. [2021], Paeckel et al. [2019], and Monte-Carlo techniques Alvarez and Gros [2002], Michel et al. [2008], Grossjohann and Brenig [2010], De Raedt et al. [2017], to name a few. Additionally, the method of dynamical quantum typicality Heitmann et al. [2020], Jin et al. [2021] is a multifaceted method for investigating transport in closed quantum systems. The basic idea of dynamical quantum typicality is that one can imitate the expectation value of an entire ensemble by an expectation value of a single pure state, drawn at random from a high-dimensional Hilbert space. Dynamical quantum typicality has been used for linear-response functions, mostly at high temperatures Bertini et al. [2021]. For such temperatures, the statistical error is smaller than at low temperatures. Still, dynamical quantum typicality is a priori neither restricted to linear-response functions nor high temperatures.
@@
In this work, we demonstrate the usefulness of dynamical quantum typicality in a specific physical situation, which differs from most previous applications in two ways. First, we apply it to energy flow in bipartite systems. Second, we also explore low temperatures. We do so for various spin-1/2 chains, which all feature ballistic energy transport: the XX chain, the critical transverse-field Ising chain, and the XXZ chain. We compare our numerical results to existing analytical results from conformal field theory (CFT) Bernard and Doyon [2012], Fischer et al. [2020] and generalized hydrodynamics (GHD) Bertini et al. [2016], Castro-Alvaredo et al. [2016], Doyon et al. [2025]. We find convincing agreement in all considered subsystem configurations of our bipartite setup, as well as for all examined temperatures. This agreement is not only found for the contact current but also for all local currents and densities in the steady state.
@@
Our paper is structured as follows: In Sec. II, we define the bipartite setup used throughout the paper and the energy current between the two subsystems. After that, we introduce the different spin-$1/2$ chains in Sec. III, which we use for our subsystem configurations. Next, in Sec. IV, we discuss our quantum typicality approach to the energy current and briefly review the CFT results. Thereafter, we present our numerical results and compare them to analytical results in Sec. V. We conclude in Sec. VI and give directions for further research.
@@ II Setup
In our paper, we are interested in the heat current from the left half into the right one. This heat current is given by Giazotto et al. [2006]
@@
Figure 2: Time evolution of the energy current for temperatures $T_{L}/J=1/6,T_{R}/J=1/8$, total system size $L=28$ for the coupling of (a) two XX chains, (b) two critical Ising chains and (c) one XX chain to a critical Ising chain. Numerical data for two different pure random initial states (NN=1) and for the average over many initial states (N=100) are compared to the CFT value. Note that larger fluctuations at longer times do not contradict the bound in Eq. (20), since this bound is not tight.
@@ III Models
First, we introduce the spin-$1/2$ XXZ model. For this model, the local Hamiltonians are given by Bertini et al. [2021]
@@
where $S_{r}^{i}\ (i=x,y,z)$ are the spin-$1/2$ operators at site $r$, $J>0$ is the antiferromagnetic exchange coupling constant, and $\Delta$ is the anisotropy in $z$ direction. Using the Jordan-Wigner transformation Jordan and Wigner [1928], the spin-$1/2$ XXZ chain can be mapped onto a chain of spinless fermions. After this transformation, the local Hamiltonians in Eq. (6) take on the form
$h_{\text{XXZ},r}=$ $\frac{J}{2}\big(c_{r}^{\dagger}c_{r+1}+\text{h.c.}\big)$
@@
Figure 3: Similar data as the one in Fig. 2 but now for different total system sizes $L$. Here, the numerical data are again averaged over $N=100$ pure random initial states. For larger system sizes $L$, finite size effects are visible for later times tJtJ.
The second model we consider is the special case of the spin-$1/2$ XXZ chain with $\Delta=0$, i.e., the XX chain. The local Hamiltonian reads
@@
where the constant $h$ denotes the strength of the transverse magnetic field. Due to the Jordan-Wigner transformation, the transverse-field Ising model is, up to a constant, equivalent to the Kitaev chain Lieb et al. [1961], Pfeuty [1970], Kitaev [2001], Suzuki et al. [2013],
@@
In our paper, we set $J_{C}=1$ for all configurations. All chosen coupling Hamiltonians ensure that energy can flow between both subsystems. Before we proceed to our numerical results for the energy current in the various subsystem configurations, we introduce our dynamical typicality approach. We also briefly discuss the result obtained within CFT Bernard and Doyon [2012].
@@ IV.1 Dynamical typicality
We now turn to a discussion of dynamical quantum typicality (DQT) Jin et al. [2021], Heitmann et al. [2020], which is the key concept used in our numerical approach. We first focus on the homogeneous case of a single Hamiltonian $H$ and a single temperature $T$, and afterwards come to the bipartite case with two $H_{L}$, $H_{R}$ and two $T_{L}$, $T_{R}$.
The basic idea of DQT Gemmer and Mahler [2003], Goldstein et al. [2006], Popescu et al. [2006] is that one can approximate the expectation value of an ensemble density matrix,
@@
where $O$ is a local operator. For such a local operator, the moments
$M_{n}=\frac{\text{tr}[O^{n}]}{D}$ (16)
with the Hilbert-space dimension $D$, do not depend on system size $L$. Moreover, we assume that these moments are finite.
If $\rho$ was the microcanonical ensemble and $|\psi\rangle$ an energy eigenstate, this idea would be identical to the well-known eigenstate thermalization hypothesis (ETH) Deutsch [1991], Srednicki [1994], Rigol et al. [2008]. In contrast to ETH, however, DQT is no assumption and relies on a particular construction of the pure state $|\psi\rangle$. Specifically, for a given ensemble density matrix $\rho$, this construction reads
@@
which is the action of the square-root of $\rho$ on a so-called Haar-random pure state $|\Phi\rangle$ Bartsch and Gemmer [2009]. As discussed below, this construction particularly enables the reduction of memory in numerical simulations, since pure states (of dimension $D$) instead of density matrices (of dimension $D^{2}$) can be used.
The particular pure state $|\Phi\rangle$ does neither depend on $\rho$ nor $H$. In an arbitrarily chosen orthogonal basis, it is given by
@@
where $D$ is again the Hilbert-space dimension and $c_{m}$ are complex coefficients. Both, the real and imaginary part of these coefficients are drawn at random according to a Gaussian probability distribution with zero mean. For such a $|\Phi\rangle$ and the resulting $|\psi\rangle$, DQT then leads to the approximation Reimann [2007], Bartsch and Gemmer [2009]
@@
where the error $\varepsilon$ depends on the particular realization of the random state $|\Phi\rangle$. However, it possible to make a statistical statement and $\varepsilon$ vanishes on average. More importantly, however, the standard deviation is bounded from above,
$\sigma(\varepsilon)\leq b={\cal O}\left(\frac{1}{\sqrt{D_{\text{eff}}}}\right)\,.$ (20)
Here, $D_{\text{eff}}$ is the so-called effective dimension. In the case of the canonical ensemble $\rho\propto\exp(-\beta H)$, which is most relevant in the context of our work and explained in more detail in Appendix A, the effective dimension is given by the partition function
@@
with the ground-state energy $E_{0}$. Thus, $D_{\text{eff}}$ counts the number of thermally occupied energy eigenstates. In the high-temperature limit $T\to\infty$, $D_{\text{eff}}=D=2^{L}$ and hence the statistical error $\varepsilon$ decreases exponentially fast with system size $L$. In the zero-temperature limit $T\to 0$, however, $D_{\text{eff}}=1$ and thus the statistical error $\varepsilon$ can be in principle large. Still, one has to take into account that the bound in Eq. (20) is not tight. In fact, at $T=0$, it turns out that the statistical error becomes $\varepsilon=0$. This fact is a kind of trivial consequence for a Haar-random pure state $|\Phi\rangle$: It is also random in the energy eigenbasis and the coefficient $c_{0}$ for the ground state $|0\rangle$ might be small but likely nonzero. At $T=0$, only this coefficient survives and one has $|\psi\rangle=|0\rangle$ Balz et al. [2018]. Note that this fact is important, since we have $D_{\text{eff}}$ close to 1 for the low temperatures considered later, in contrast to the large $D_{\text{eff}}$ in usual high-temperature consideratios, e.g., $D_{\text{eff}}=D=2^{30}$ for $L=30$.
@@
if not stated otherwise. In this way, the statistical error in Eq. (20) is multiplied by the factor $1/\sqrt{L}$ Jaklič and Prelovšek [2000]. In our case, the local observable $O$ is mostly the contact current $J_{0}$, but we also consider currents $J_{r}$ at other sites $r$ as well as the local energy densities $h_{r}$.
In our numerical simulations, we calculate the action of the imaginary-time operators $e^{-\beta_{L}H_{L}/2}$ and $e^{-\beta_{R}H_{R}/2}$ on the Haar-random pure states $|\Phi_{L}\rangle$ and $|\Phi_{R}\rangle$ with a fourth-order Runge-Kutta scheme, where we choose a very small step $\delta\beta J=0.0001$. This scheme is also used for the action of the real-time operator $e^{-iHt}$ on the pure state $|\psi\rangle$, where we choose a larger but still small step $\delta tJ=0.01$. By doing so and also using sparse-matrix techniques, we can treat system sizes outside the range of standard exact diagonalization. Alternatively, other propagation methods, such as Trotter decompositions Suzuki [1985], De Raedt [1987], Berry et al. [2006] or Chebyshev polynomial expansions Tal-Ezer and Kosloff [1984], Weiße et al. [2006], can be used as well.
@@ IV.2 Conformal field theory
Figure 5: Steady-state profile for (a) the local density $\langle h_{r}\rangle$ and (b) the local current $\langle J_{r}\rangle$ for the coupling of two XXZ chains with the same anisotropies $\Delta$, $T_{L}/J=1$$T_{R}/J=1/2$, and total system size $L=24$. (Note that in the $x$-axis the position $r$ is divided by the time $t$.) The numerical data are compared to the results from GHD in Ref. Bertini et al. [2016]. Here, the numerical data are shown for both, a random product state (open symbols) and a random state in the entire Hilbert space (closed symbols).
The energy transport in a bipartite CFT setup has been studied by Bernard and Doyon Bernard and Doyon [2012]. Specifically they considered a CFT separated into two halves prepared at different temperatures. These halves were then glued together and the non-equilibrium steady state emerging at late times was studied. In particular, for the energy current they obtained the universal formula
@@
where $c$ denotes the central charge of the CFT. Given that CFTs generically describe the universal low-energy behavior of certain lattice models, the result (25) is expected to hold only in the low-temperature regime. The central charge $c$ is a specific characteristic of the CFT Di Francesco et al. [1997], physically it can be interpreted as the number of degrees of freedom, which here contribute to the energy transport through the system.
The setup leading to the universal CFT result (25) possesses translation invariance after the gluing quench. As an extension, Ref. Fischer et al. [2020] studied the coupling of an XX chain to a critical Ising chain, i.e., our setup (iii) above. Using both the non-equilibrium Green function formalism and the density matrix renormalization group method, the energy current was found to behave as
@@
This result can be viewed as a bottleneck effect, where the subsystem with the lower number of degrees of freedom as reflected by the lower central charge limits the energy transport. The result (26) is also consistent with a purely field theoretic analysis by Bernard et al. Bernard et al. [2015], where, however, the precise link to a microscopic setup remains elusive.